Petrus Framing's cost formula for its supplies cost is $1,920 per month plus $11 per frame. For the month of March, the company planned for activity of 632 frames, but the actual level of activity was 639 frames. The actual supplies cost for the month was $9,340. The activity variance for supplies cost in March would be closest to:
Answer #1

An activity variance is the difference between a revenue or cost item in the flexible budget and the same item in static planning budget.

Estimated supplies cost for 632 frames = 1,920 + 639*11= $8,949

Actual supply cost = $9,340

Total Activity variance= Actual- estimated

Total variance= 9,340 - 8,949=$ 391 unfavorable